### v2

# Selective Transformer for Hyperspectral Image Classification

Abstract
Transformer has achieved satisfactory results in the field of hyperspectral image (HSI) classification. However, existing Transformer models face two key challenges when dealing with HSI scenes characterized by diverse land cover types and rich spectral information: (1) fixed receptive field representation overlooks effective contextual information; (2) redundant self-attention feature representation. To address these limitations, we propose a novel Selective Transformer (SFormer) for HSI classification. The SFormer is designed to dynamically select receptive fields for capturing both spatial and spectral contextual information, while mitigating the impact of redundant data by prioritizing the most relevant features. This enables a highly accurate classification of the land covers of the HSI. Specifically, a Kernel Selective Transformer Block (KSTB) is first utilized to dynamically select an appropriate receptive field range to effectively extract spatial-spectral features. Furthermore, to capture the most crucial tokens, a Token Selective Transformer Block (TSTB) is introduced, which selects the most relevant tokens based on the ranking of attention scores for each query. Extensive experiments on four benchmark HSI datasets demonstrate that the proposed SFormer outperforms the state-of-the-art HSI classification models. The codes will be released.
Index Terms:
Hyperspectral image (HSI) classification, transformer, selective spatial attention, fine-grained spectral attention.
I Introduction
With the continuous advancements in hyperspectral imaging technology, hyperspectral images (HSIs) now offer increasingly rich spatial-spectral information, enabling precise Earth observation[1, 2, 3, 4]. Leveraging their distinctive spectral characteristics, HSIs are capable of performing a wide range of tasks, including classification[5, 6, 7], target detection[8, 9], change detection[10, 11], and image quality enhancement[12, 13]. Notably, HSI classification has become a fundamental task in the remote sensing community due to its extensive applications in areas such as disaster monitoring[14], precision agriculture[15], and urban planning[16]. The primary goal of HSI classification is to assign specific class labels to each pixel within the HSI.
Deep neural networks have become dominant in the field of HSI classification, including stacked auto-encoders (SAEs)[17], deep belief networks (DBNs)[18], recurrent neural networks (RNNs)[19], graph convolutional networks (GCNs)[20], convolutional neural networks (CNNs)[21], vision transformers (ViTs)[22] and recently emerging state space models (SSMs)[23]. Deep learning methods offer the advantage of automatically learning discriminative deep features from input images without relying on hand-crafted features.
Owing to their powerful feature representation capabilities, CNNs[24, 25, 26] have attracted significant attention. A 1D-CNN[27] was proposed to achieve efficient local spectral feature extraction through multiple convolutional layers. Further, Zhao et al. [28] introduced 2D-CNN that extracts spatial-spectral features and uses a balanced local discriminant embedding. Later, 3D-CNNs [29, 30] were developed to directly extract joint spatial-spectral features from the HSI cube. Then, some multiscale convolutional networks[31, 32] were proposed to extract deep multiscale features. However, these CNN-based networks are constrained by their inherent local perception, which limits the amount of information they can capture, and thus hinders accuracy.
Subsequently, HSI classification approaches with attention mechanism[33, 34, 35, 36] emerged and achieved significant classification results, particularly self-attention mechanisms[37, 38, 39], due to their ability to improve long-range perception. After that, transformer-based networks[40, 41, 42, 22, 43, 44, 45, 46] have been widely adopted for their use of Multi-Head Self-Attention (MHSA) to achieve superior context modeling. These approaches have notably outperformed traditional CNNs, which are limited by their receptive fields. Hong et al. [22] proposed SpectralFormer, a network that enhances HSI classification by capturing spectral sequence information between adjacent HSI bands using transformers and cross-layer skip connections. Further, Sun et al. [43] proposed a spectral–spatial feature tokenization transformer (SSFTT), where shallow features are extracted by convolutions, semantic features are processed by a Gaussian weighted tokenization module, and a single-layer transformer is used for feature representation. Then, Zou et al. [47] introduced LESSFormer, which employs an HSI2Token module to extract adaptive spatial-spectral tokens and uses a local transformer to enhance the local spectral representation. Inspired by the effectiveness of spectral partitioning, Mei et al. [48] introduced a grouped pixel embedding strategy and hierarchical structure to extract discriminative multiscale spatial-spectral features. Roy et al. [49] incorporated morphological characteristics into their proposed MorphFormer by combining spectral and spatial morphological convolution operations with attention mechanisms to effectively integrate structural and morphological information. After that, Zhao et al. [45] introduced a lightweight variant of the ViT network, which combines group-separable convolution and Multi-Head Self-Attention modules within feature extraction blocks to efficiently capture local and global spatial features with reduced parameter complexity.
However, most transformer structures typically utilize a fixed receptive field for feature extraction, overlooking the fact that different types of land cover require varying contextual information[50, 51]. As shown in Fig. 1 (a), HSI with a limited receptive field may exhibit misclassification issues. Different types of land cover often require varying ranges of contextual information to allow an accurate classification of pixels within the HSI. Moreover, most transformer-based HSI classification models usually use the vanilla Multi-Head Self-Attention to build global features[48]. Fig. 1 (b) illustrates tokens of the land cover category related to the query token, as well as some tokens not related to the type of land cover target. In addition, existing transformer-based HSI classification models predominantly focus on spatial attention or spectral attention[22], neglecting the integrated modeling of spatial and spectral contextual information within HSI data.
To address these issues, this paper proposes a novel Selective Transformer (SFormer) for HSI classification, which selects the most valuable and relevant spatial-spectral information within a reasonable receptive field. SFormer is designed with the understanding that objects of different scales require varying contextual information. To accommodate this, the model incorporates a Kernel Selection Transformer Block (KSTB), which dynamically selects kernels of appropriate sizes, optimizing the receptive fields for effective spatial-spectral feature extraction. Furthermore, rather than utilizing all tokens for dense attention calculation, SFormer introduces a Token Selective Transformer Block (TSTB). This block selects the most relevant tokens with the highest attention values with a group strategy that can directly utilize the 3D HSI characteristic. These selective mechanisms allow SFormer to capture the most significant and appropriate spatial-spectral information within an optimal receptive field. The main contributions of this study are summarized as follows:
Figure 1: (a) An illustration of limited receptive fields in HSI classification shows how varying land-cover types require different scales of context, potentially leading to misclassification, such as confusing grass with bricks. (b) An example of relevant and irrelevant tokens in relation to a query token demonstrates that while relevant tokens provide valuable information for accurately classifying an asphalt land-cover type, irrelevant tokens may introduce noise, potentially impairing classification accuracy.
  • 1)
    This paper proposes a novel Selective Transformer (SFormer) for HSI classification. The SFormer effectively selects receptive fields to extract both spatial and spectral contextual information, while reducing redundant information interference by focusing on the most relevant features. This approach enables the precise classification of HSI ground objects.
  • 2)
    To adaptively select the appropriate contextual information, we design a Kernel Selection Transformer Block (KSTB). This mechanism ensures that feature extraction occurs within an optimal receptive field from both spatial and spectral perspectives, thereby enhancing the model’s capacity to effectively learn and distinguish diverse land cover features.
  • 3)
    We further develop a Token Selective Transformer Block (TSTB) that strategically selects the most relevant tokens to extract valuable information. This mechanism effectively integrates essential information while reducing the impact of redundant and irrelevant information, thereby improving the representation of spatial-spectral features in visual tasks.
  • 4)
    Extensive experiments are conducted on four benchmark HSI datasets: Pavia University, Houston, Indian Pines, and WHU-HongHu. The results indicate that the proposed SFormer, which utilizes appropriate and valuable contextual information, performs better than other state-of-the-art methods.
The remainder of this paper is organized as follows. Section II provides a comprehensive description of the proposed SFormer model. Section III presents and thoroughly analyzes the experimental results. Finally, Section IV concludes the paper.
II Proposed Methods
In this section, we present the implementation details of our SFormer. SFormer consists of convolutional layers, selective transformer groups (STG), and a classification head. We begin with an overview of SFormer, followed by a detailed explanation of its components: the Kernel Selective Transformer Block (KTAB), and the Token Selective Transformer Block (TSTB).
Figure 2: An illustration of the proposed SFormer. The Selective Transformer Group (STG) is composed of a Kernel Selective Transformer Block (KSTB) and three consecutive Token Selective Transformer Blocks (TSTBs). Followed by two STGs, we adopt the fully connected layer as the classification head to obtain the classification results.
II-A Overview of SFormer
Figure 3: The proposed Kernel Selective Attention (KSA) conceptual illustration. Here, DwConv represents Depth-wise convolution, Avg and Max represent channel-wise average pooling and max pooling, respectively, and GAP means spatial global average pooling.
As shown in Fig. 2, the HSI is first subjected to PCA for dimensionality reduction, followed by the extraction of overlapping patch cubes, $\{x_{i}\in\mathbb{R}^{c\times w\times w}\}_{i=1}^{H\times W}$, where each patch is classified based on the central pixel. Here, $H$ and $W$ denote the height and width, $w$ represents the patch size, and $c$ is the reduced spectral dimension. Taking the pixel-level patch cube $x\in\mathbb{R}^{c\times w\times w}$ as input, we first employ a 3×3 convolutional layer to capture shallow features, denoted as $\mathbf{F}_{1}$. Next, we perform deep feature extraction on $\mathbf{F}_{1}$ using two consecutive Selective Transformer Groups (STGs). This process can be formulated as: $\mathbf{F}_{2}=\text{STG}_{2}(\text{STG}_{1}(\mathbf{F}_{1}))$. Finally, a layer normalization followed by a fully connected layer in the classification head is applied to obtain the final classification results.
II-B Kernel Selective Transformer Block
As analyzed in the Introduction, different land cover classes in HSI require distinct contextual information. Existing methods are often limited by the fixed size of the receptive field of the network or cannot dynamically adapt the receptive field to suit specific types of land cover[48]. To address this issue, we designed the Kernel Selective Transformer Block (KSTB). First, it employs dilated depthwise convolutions to construct a larger receptive field for multiscale modeling. Second, it incorporates a spatial-spectral selection mechanism to adaptively determine the appropriate receptive field size required for different types of land cover. Specifically, KSTB primarily consists of a Kernel Selective Attention (KSA) module and a Feed-Forward Network (FFN).
The structure of KSA is shown in Fig. 3, two decoupled depth-wise convolutions (DwConv) are employed for input $\mathbf{X}\in\mathbb{R}^{c\times w\times w}$. The features generated with decomposed kernels are then processed by 1×1 convolution layers $\mathcal{F}_{1\times 1}(\cdot)$ to integrate channel information and unify dimensions, obtaining contextual features with different receptive fields:
$\mathbf{{U}_{1}}=\mathcal{F}^{1}_{1\times 1}(\mathcal{F}_{3\times 3}^{dwc}(\mathbf{X})),\mathbf{{U}_{2}}=\mathcal{F}^{2}_{1\times 1}(\mathcal{F}_{5\times 5}^{dwc}(\mathcal{F}_{3\times 3}^{dwc}(\mathbf{X}))),$ (1)
where $\mathcal{F}_{3\times 3}^{dwc}$ and $\mathcal{F}_{5\times 5}^{dwc}$ denote a 3×3 DwConv and a 5×5 DwConv with a dilation of 2, respectively.
Given the varying scales of contextual information required by different types of land cover, we employ a spatial-spectral selection mechanism to enhance the dynamic adaptability of the network. Specifically, we first concatenate the features $\mathbf{{U}_{1}},\mathbf{{U}_{2}}\in\mathbb{R}^{\frac{c}{2}\times w\times w}$ extracted by convolutional kernels with diverse receptive fields:
$\mathbf{U}=\left[\mathbf{{U}_{1}};\mathbf{{U}_{2}}\right]\in\mathbb{R}^{c\times w\times w},$ (2)
subsequently, channel-wise average pooling $\mathcal{P}_{\text{avg }}(\cdot)$ and max pooling $\mathcal{P}_{\text{max}}(\cdot)$ are applied to squeeze $\mathbf{U}$ to capture spatial descriptors. To enable information exchange among different spatial descriptors, we first concatenate the aforementioned pooling features and then using a simple 1×1 convolutional layer $\mathcal{F}^{2\rightarrow O}_{1\times 1}$ to transform the two pooling features into $O$ spatial attention maps, we set $O$ to 2 in our structure. These process can be written as below:
$\mathcal{\bar{\mathbf{S}}}=\mathcal{F}^{2\rightarrow O}_{1\times 1}(\left[\mathcal{P}_{\text{avg}}(\mathbf{U});\mathcal{P}_{\text{max}}(\mathbf{U})\right]).$ (3)
Following this, the sigmoid activation function $\sigma(\cdot)$ is applied to each spatial attention feature map $\mathcal{\bar{\mathbf{S}}}\in\mathbb{R}^{2\times w\times w}$, generating separate spatial selection masks corresponding to the different convolutional kernels:
$\mathbf{S}_{i}=\sigma(\mathcal{\bar{\mathbf{S}}}_{i}).$ (4)
Considering the substantial differences in contextual spectral information required by various types of complex land cover in HSIs, we have devised a spectral selection mechanism to meet this demand. First, we utilize spatial global average pooling $\mathcal{P}_{\text{GAP }}(\cdot)$ to compress the feature $\mathbf{U}$ into spectral feature descriptors. Next, a fully connected layer $\mathcal{F}_{\text{FC }}(\cdot)$ is employed to produce more compact spectral attention features $\boldsymbol{Z}\in\mathbb{R}^{c\times 1}$. These operations can be mathematically expressed as follows:
$\boldsymbol{Z}=\mathcal{F}_{\text{FC }}(\mathcal{P}_{\text{GAP }}(\mathbf{U})).$ (5)
Subsequently, spectral-wise attention is generated through a softmax operation to guide the selection process, resulting in spectral selection masks, which can be obtained as follows:
$C_{1}^{[j]}=\frac{e^{{M_{j}}{Z_{j}}}}{e^{{M_{j}}{Z_{j}}}+e^{{N_{j}}{Z_{j}}}},C_{2}^{[j]}=\frac{e^{{N_{j}}{Z_{j}}}}{e^{{M_{j}}{Z_{j}}}+e^{{N_{j}}{Z_{j}}}},$ (6)
where $C_{i}^{[j]}$ represents $j$-th attention value in $\boldsymbol{C_{i}}\in\mathbb{R}^{c\times 1},i=1,2$, $M_{j}$ refers the $j$-th row of $\boldsymbol{M}$, likewise $C_{2j}$ and $N_{j}$, and $\boldsymbol{M},\boldsymbol{N}\in\mathbb{R}^{c\times 1}$ are two initially defined learnable vectors.
Furthermore, the spatial-spectral selection weights are obtained by matrix multiplication with the corresponding spatial-spectral selection masks:
$\mathbf{W}_{1}=\boldsymbol{C}_{1}\times\mathbf{S}_{1},\mathbf{W}_{2}=\boldsymbol{C}_{2}\times\mathbf{S}_{2}.$ (7)
The enhanced feature maps derived from the decomposed kernel sequence are weighted using their respective spatial-spectral selection weights $\mathbf{W}_{i}\in\mathbb{R}^{c\times w\times w}$. These weighted maps are then combined via the convolutional layer $\mathcal{F}_{1\times 1}(\cdot),i=1,2$, the attention feature $\mathbf{S}$ can be expressed as:
$\mathbf{S}=\mathcal{F}_{1\times 1}(\sum_{i=1}^{O}\left(\mathbf{W}_{i}\cdot\mathbf{U}_{i}\right)).$ (8)
The KSA module generates its final output by performing an element-wise multiplication between the input features $\mathbf{X}$ and $\mathbf{S}$:
$\mathbf{\mathbf{{X}^{{}^{\prime}}}}=\mathbf{X}\cdot\mathbf{S}.$ (9)
The FFN module comprises a fully connected layer, a depthwise convolution, a GELU activation function, and a second fully connected layer.
II-C Token Selective Transformer Block
Previous research has typically utilized the self-attention mechanism within Transformers to capture long-range dependencies, facilitating the modeling of global spatial or spectral relationships[22]. However, HSI often contains a wide variety of complex land cover types. The conventional Transformer architecture tends to compute attention across all query-key pairs, which inevitably leads to interactions between redundant and irrelevant information, ultimately diminishing interpretative accuracy.
Figure 4: The illustration of the proposed Multi-Head Selective Attention (MSA). The input feature $P$ is first grouped and then fed through a 3D Point-wise convolution, followed by a 3D Depth-wise convolution to generate the corresponding $\mathbf{Q}\mathbf{K}\mathbf{V}$. Subsequently, a Token selection mechanism is utilized for further self-attention operation.
To overcome this limitation, we propose a novel multi-head selective attention (MSA) mechanism to replace the traditional multi-head self-attention (MHSA) mechanism. Furthermore, our MSA retains the data characteristics of the HSI three-dimensional cube through a grouped operation and employs 3D convolutions to extract tokens with spatial and spectral properties. During the computation of self-attention, the MSA selectively focuses on the most relevant tokens, effectively disregarding the least relevant information. This approach improves interpretive accuracy while reducing redundant computations.
As illustrated in Fig. 4, we first begin by partitioning the input feature $\mathbf{P}\in\mathbb{R}^{C\times H\times W}$ into $g$ groups, each producing a feature map of size $c\times H\times W$. Then we feed them through a 3D point-wise convolution with a kernel size of $1×1×$ 1 × 1, followed by a 3D depth-wise convolution with a kernel size of $1×3×3$. This process generates the corresponding query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$. To facilitate the calculation of self-attention, we reshape and transpose the tensors $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, adjusting their dimensions by flattening the operation, so each has a shape of $(H\times W\times g)\times c$. These processes can be extended to the multi-head scenario and expressed as follows:
$\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}=Split(\mathcal{F}_{1\times 3\times 3}^{dwc}(\mathcal{F}_{1\times 1\times 1}(\mathbf{e}_{i}))),$ (10)
$\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}=Flatten(Reshape(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})),$ (11)
where $\mathbf{e}_{i}$ denotes the feature in $i$th head, $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{HWg\times c}$ correspond to $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ in $i$th head, respectively, and $g$ represents the number of groups. The $Split(\cdot)$ function refers to the chunk operation performed along the channel dimension. Next, a dense attention matrix $\mathcal{A}\in\mathbb{R}^{HWg\times HWg}$ is generated through the dot product between $\mathbf{Q}_{i}$ and $\mathbf{K}_{i}$:
$\mathcal{A}=\frac{{\mathbf{Q}_{i}}{\mathbf{K}_{i}}^{T}}{\lambda},$ (12)
where $\lambda=\sqrt{\frac{C}{h}}$, $h$ denotes the number of heads. Following this, the token selection mechanism is applied to identify the top $k$ percent of the most relevant elements within $\mathcal{A}$. For example, with $k=0.8$, the top 80% of the elements are retained for activation, while the remaining 20% are masked to 0. To achieve this, a matrix $\mathcal{M}^{k}$ is created, which operates as follows:
$\mathcal{M}^{k}_{pq}=\begin{cases}\mathcal{A}_{pq},&\mathcal{A}_{pq}>T_{k}^{[p]}\\ -\infty,&\text{otherwise},\end{cases}$ (13)
where $\mathcal{M}^{k}_{pq}$ denotes the value at position $pq$, and $\boldsymbol{T_{k}}\in\mathbb{R}^{HWg\times 1}$ is the lower threshold vector of the highest $k$ percent attention values in each row of $\mathcal{A}$, i.e, each token has a unique threshold, and $T_{k}^{[p]}$ is the $p$th value of $\boldsymbol{T_{k}}$. After that, the selective attention matrix is activated by the softmax function. The activated attention can be written as follows:
Algorithm 1 Token Selective Attention
Input: Feature $\mathbf{P}\in\mathbb{R}^{C\times H\times W}$, group number $g$, head number $h$, token selection rate $k$.
Initialization: $\mathbf{e}_{i}\in\mathbb{R}^{d\times H\times W}$ is the $i$-th head of feature, $d=\frac{C}{h}$. $\mathcal{M}^{k}\in\mathbb{R}^{n\times n}$ is initialized with $\infty$, $n=HWg$. $c=\frac{d}{g}$, $\lambda=\sqrt{d}$.
1:  for each $\mathbf{e}_{i}$ do
2:     $\mathbf{G}_{i}\leftarrow Group(\mathbf{e}_{i})$; // $g\times c\times H\times W$
3:     $\mathbf{G}_{i}\leftarrow DWConv(PConv(\mathbf{G}_{i}))$; // $g\times c\times H\times W$
4:     $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\leftarrow Flatten(Chunk(\mathbf{G}_{i}))$; // $HWg\times c$
5:     $\mathcal{A}\leftarrow\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\lambda}$; // Dense Attention $HWg\times HWg$
6:     $\boldsymbol{T}_{k}^{[p]}\leftarrow\min(Topk\{\mathcal{A}_{p1},\cdots,\mathcal{A}_{pn}\},p=1,\cdots,n)$;// $1\times 1$
7:     $\mathcal{M}_{pq}^{k}\leftarrow\mathcal{A}_{pq}\;if\;\mathcal{A}_{pq}>T_{k}^{[p]}$; // $HWg\times HWg$
8:     $\mathcal{A}^{*}\leftarrow Softmax(\mathcal{M}^{k})$; // Selective Attention $HWg\times HWg$
9:     $\mathbf{V}^{*}_{i}\leftarrow\mathcal{A}^{*}\otimes\mathbf{V}_{i}$; // $HWg\times c$
10:     $\mathbf{V}^{*}_{i}\leftarrow Reshape(\mathbf{V}^{*}_{i})$; // $d\times H\times W$
11:     $\mathbf{P}^{{}^{\prime}}_{i}\leftarrow\mathbf{P}_{i}+\mathbf{V}^{*}_{i}$ // $d\times H\times W$
12:  end for
13:  $\mathbf{P}^{{}^{\prime}}\leftarrow\mathcal{F}_{1\times 1}(Concat(\mathbf{P}^{{}^{\prime}}_{i}))$ // $C\times H\times W$
$\mathcal{A}^{*}=Softmax(\mathcal{M}^{k}),$ (14)
Further, the $i$th head selective attention $\mathbf{V}^{*}_{i}$ can be multiplied by $\mathcal{A}^{*}$ and $\mathbf{V}_{i}$, and then add the input feature $\mathbf{P}_{i}$ for residual connection:
$\mathbf{P}^{{}^{\prime}}_{i}=\mathbf{P}_{i}+\mathcal{A}_{i}^{*}\mathbf{V}_{i}.$ (15)
Given our multi-head design, we concatenate the outputs from each head and then aggregate them using a 1×1 convolution, as demonstrated below:
$\mathbf{P^{{}^{\prime}}}=\mathcal{F}_{1\times 1}(Concat(\mathbf{P}^{{}^{\prime}}_{i})).$ (16)
The detailed implementation of our Token Selective Attention is presented in Algorithm 1.
III Experiments
In this section, we first provide an overview of the datasets and the implementation details. We then conduct a comprehensive evaluation of the proposed method, including in-depth parameter analysis and module ablation studies. This is followed by a qualitative and quantitative comparison of our approach with current state-of-the-art methods. Finally, we evaluated and discussed the operational efficiency of our model compared to other methods.
Figure 5: Comparison of classification performance across four datasets under different values of $k$.
Figure 6: Comparison of classification performance across four datasets under different group number $g$.
III-A Data Descriptions
III-A1 Pavia University
This dataset was acquired by the Reflective Optics System Imaging Spectrometer (ROSIS) on the campus of Pavia University in northern Italy in 2001. It has 103 spectral bands. The dataset consists of 610 by 340 pixels and covers 9 main categories of interest.
III-A2 Houston
The Houston dataset was acquired over the University of Houston campus and its surrounding areas in 2012[52]. It has 144 spectral bands in the wavelength range of 400-1000 nm and a spatial resolution of 2.5 m. It consists of 349 by 1905 pixels and covers 15 categories.
III-A3 Indian Pines
The AVIRIS sensor acquired this dataset in North-west Indiana in 1992. It has a dimension of 145 by 145 pixels and a spatial resolution of 20m. It covers 200 spectral bands in the wavelength range of 400-2500 nm after removing the water absorption bands. The dataset includes 16 vegetation classes.
III-A4 Whu-HongHu
This dataset was collected using a UAV platform in HongHu City in 2017[53]. It comprises 270 spectral bands ranging from 400 to 1000 nm, with a high spatial resolution of approximately 0.04m. The dimensions of the imagery are 940 by 475 pixels, which includes 17 typical crop types.
III-B Experimental Settings
The experiments in this study were conducted on the PyTorch platform, utilizing an NVIDIA GeForce RTX 3090 GPU. For the Pavia University dataset, 30 samples per class were randomly selected for the training set, while for the other three datasets, 50 samples per class were chosen, with the remaining labeled samples in all cases used as the test set. The training process used the AdamW optimizer, starting with an initial learning rate of $41e-$ 4 and a weight decay of $51e-$ 5, was run for 500 epochs, extract patch cubes of size $C\times 10\times 10$, patch size in transformer is 2, and embedding dimension is fixed with 128 in all datasets. To quantitatively assess classification performance, we used three widely recognized evaluation metrics: overall accuracy (OA), average accuracy (AA), and kappa coefficient ($\kappa$). Each experiment was repeated 10 times with different random seeds, and the classification results are presented as mean and standard deviation.
TABLE I: Performance Contribution of Each Module on four datasets (reported in OA). Best Results Are Highlighted in Bold. All results are in %.
Depth KSTB TSTB Pavia University Houston Indian Pines HongHu
1 87.67 92.51 85.97 88.32
9 86.33 90.04 85.99 86.58
9 93.44 96.98 92.14 92.95
9 95.81 97.44 93.65 93.84
9 96.59 97.66 95.17 94.59
TABLE II: Exploring the impact of Spatial-Spectral Selection mechanism on performance on four datasets (reported in OA). The best results are highlighted in Bold. All results are in %.
SFormer Pavia University Houston Indian Pines HongHu
w/o Spe Selection 96.09 97.43 94.38 94.32
w/o Spa Selection 96.01 97.41 94.20 94.13
w/ Spa-Spe Selection 96.59 97.66 95.17 94.59
“w/o Spa/Spe Selection” and “w/ Spa-Spe Selection” represent without Spatial/Spectral Selection and Spatial-Spectral Selection, respectively.
III-C Parameter Analysis
TABLE III: Quantitative Classification Results of the Pavia University Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
1 77.43±7.99 77.51±1.48 80.10±3.43 82.17±3.72 66.00±11.47 77.11±17.13 88.39±9.29 87.79±6.34 93.48±3.62 96.08±1.65
2 70.94±10.10 78.88±2.21 80.12±3.82 90.74±2.95 77.25±8.09 69.92±20.30 89.27±4.43 90.40±4.98 91.69±6.45 95.76±3.15
3 68.47±14.58 76.25±2.10 82.13±4.84 95.35±2.63 80.32±11.09 60.63±14.57 87.25±6.23 74.95±16.35 94.09±4.18 95.57±2.45
4 93.48±7.12 89.31±1.52 77.97±3.18 86.59±4.38 91.10±2.12 97.26±2.03 96.87±1.65 93.50±3.43 93.61±4.88 97.09±0.76
5 99.56±0.42 98.09±1.53 95.70±1.47 99.64±0.36 99.86±0.21 99.85±0.22 99.85±0.19 99.54±0.51 99.86±0.15 99.94±0.12
6 75.85±13.08 82.65±2.37 91.96±2.09 96.54±2.14 77.09±11.96 77.10±16.86 92.96±5.98 94.20±4.21 93.86±4.74 98.00±1.35
7 88.00±3.75 86.72±1.84 92.35±2.37 95.45±0.89 75.66±10.84 78.88±20.94 97.07±3.82 93.50±5.26 98.53±1.17 99.51±0.42
8 82.22±15.26 84.50±2.23 86.65±1.12 90.72±3.21 62.86±20.38 76.74±20.03 88.08±5.43 91.89±6.18 95.12±1.94 97.04±0.85
9 99.42±0.67 99.09±0.63 98.46±0.77 95.53±4.20 96.80±1.23 98.83±1.95 99.35±0.49 99.75±0.25 99.64±0.44 99.12±0.66
OA (%) 77.01±4.48 81.48±1.45 83.27±1.72 90.54±2.17 76.46±2.36 75.79±9.10 90.69±2.44 90.62±2.29 93.40±2.43 96.59±1.41
$\kappa$ (%) 70.95±5.22 76.23±1.78 78.53±2.03 87.68±2.74 69.94±2.64 69.92±9.98 87.91±3.09 87.77±2.89 91.39±3.05 95.52±1.82
AA (%) 83.93±2.36 85.89±0.98 87.27±0.91 92.53±1.66 80.77±1.91 81.81±4.10 93.23±1.38 91.72±1.69 95.54±1.07 97.57±0.51
Figure 7: Visualization of the classification maps on the Pavia University dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-C1 Effect of Different Values $k$
In the TSTB module, one of the core parameters to select the $k$ tokens is the value of $k$. To identify the optimal value of $k$, we performed four experiments with $k$ set to 0.2, 0.4, 0.6, 0.8, and 1, respectively. When $k$ is set to 0.8, the TSTB module selects the most relevant 80% tokens for the spatial-spectral context modeling, while the remaining 20% are considered the least relevant and are therefore excluded from the information interaction process. When $k$ is set to 1, the attention mechanism in the TSTB module degenerates into the conventional full self-attention form. Fig. 5 presents the OA results in four datasets with different values of $k$. The experimental results show that when $k$ is set to 0.2, the classification accuracy is relatively low due to the insufficient information available for long-range modeling. As the value of $k$ increases, the OA improves, indicating that more pixels contribute to the construction of contextual relationships, enabling the network to capture richer and more effective information for the classification task. However, when $k$ exceeds the optimal value, OA begins to decline, likely due to the inclusion of a significant amount of irrelevant or redundant information in the remaining $1-k$ tokens, which negatively impacts classification performance. Finally, based on the experimental results, we selected $k$ values of 0.4, 0.6, 0.8, and 0.8 across the four datasets to achieve optimal classification performance. Furthermore, performance in most $k$ configurations outperforms the traditional self-attention mechanism, validating the effectiveness of the token selection strategy.
III-C2 Effect of Different Group Numbers $g$
A critical parameter in the SFormer architecture is the number of groups, denoted as $g$. To determine the optimal value for $g$, we conducted four experimental configurations with $g$ set to 1, 2, 4, and 8, respectively. By utilizing grouping and 3D convolutions, the TSTB module preserves the spatial-spectral characteristics of HSIs, enabling the extraction of tokens that incorporate both spatial and spectral features. The value of $g$ directly impacts the size of the attention matrix, with larger $g$ values leading to increased computational complexity due to the expansion of the matrix. When $g$ is set to 1, the TSTB module reduces to purely spatial token selection without any grouping. The experimental results, as illustrated in Fig. 6, show that when $g$ is set to 1, no grouping strategy is applied to preserve spatial-spectral characteristics, resulting in a relatively lower classification accuracy across the four datasets. In contrast, the grouping strategy significantly improves classification performance. Notably, the optimal classification performance is achieved with $g$ is set to 4 across all four datasets, while maintaining relatively efficient computational complexity.
TABLE IV: Quantitative Classification Results of the Houston Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
1 91.96±6.02 94.17 ± 1.94 87.38 ± 1.43 94.44 ± 1.57 94.32±3.37 87.76±5.05 93.61±4.89 96.79±1.73 96.81±1.95 98.25±1.26
2 92.57±6.08 96.80 ± 2.54 79.50 ± 3.08 90.35 ± 1.89 96.45±1.04 94.97±5.43 97.52±3.83 97.57±3.50 99.31±0.86 99.37±0.34
3 99.21±0.95 99.09 ± 0.49 98.56 ± 0.87 99.41 ± 0.64 99.43±0.42 90.36±17.37 99.57±0.49 99.32±0.90 99.58±0.54 99.61±0.26
4 93.58±2.12 89.61 ± 1.12 82.05 ± 3.02 90.59 ± 2.89 93.72±2.53 97.85±2.74 96.81±2.57 96.73±2.60 99.39±0.65 99.56±0.67
5 98.98±0.83 97.23 ± 0.58 82.10 ± 3.33 97.87 ± 0.80 98.70±0.92 98.72±1.85 99.66±0.57 99.35±0.82 100 99.87±0.24
6 92.95±5.28 91.78 ± 1.09 89.20 ± 2.17 95.85 ± 1.68 96.33±1.87 97.09±3.11 98.95±1.18 98.87±1.19 99.96±0.11 99.78±0.65
7 87.83±2.82 82.31 ± 1.36 73.04 ± 1.83 85.48 ± 2.66 78.10±4.61 87.73±6.38 94.11±2.77 93.45±2.06 95.48±2.67 97.95±1.06
8 72.97±6.39 83.15 ± 2.30 62.19 ± 2.66 80.45 ± 3.09 84.86±3.65 78.49±5.71 89.36±2.98 87.24±4.22 89.66±4.75 92.28±2.94
9 83.60±4.00 81.41 ± 1.20 68.82 ± 2.52 80.05 ± 2.10 76.15±4.90 89.54±5.53 90.87±3.73 90.52±2.48 90.87±4.30 94.38±2.07
10 85.53±5.68 88.07 ± 1.97 90.93 ± 2.32 92.56 ± 3.15 89.97±3.90 63.59±27.82 93.02±5.28 96.29±2.30 98.09±1.25 97.69±3.11
11 71.20±4.31 84.08 ± 2.49 84.14 ± 1.93 94.36 ± 2.81 76.36±3.41 82.01±14.44 94.18±2.62 93.79±2.55 96.25±2.02 98.03±1.81
12 84.33±4.93 79.37 ± 2.80 84.57 ± 2.84 94.61 ± 1.99 80.35±4.81 67.31±29.28 92.53±3.96 94.22±3.13 93.44±4.63 95.16±2.97
13 94.84±0.66 88.76 ± 1.74 82.98 ± 3.61 97.28 ± 1.33 76.11±4.69 96.56±2.74 97.18±1.48 98.42±1.60 98.59±1.65 99.38±0.66
14 98.68±1.31 98.39 ± 0.48 98.92 ± 0.69 100 96.85±1.76 99.79±0.39 100 99.81±0.56 100 99.97±0.08
15 99.64±0.39 95.77 ± 0.60 94.23 ± 2.15 99.64 ± 0.32 98.84±1.08 99.82±0.21 99.89±0.18 99.69±0.42 99.66±0.42 100
OA (%) 88.14±1.32 88.89±1.00 81.74±0.93 91.47±1.32 88.10±0.74 86.72±4.09 94.99±0.91 95.36±0.99 96.52±0.60 97.66±0.51
$\kappa$ (%) 87.17±1.43 87.99±1.08 80.26±1.00 90.78±1.42 87.14±0.80 85.65±4.41 94.58±0.99 94.98±1.07 96.23±0.65 97.47±0.56
AA (%) 89.86±1.18 90.00±0.85 83.91±0.90 92.86±1.07 89.10±0.68 88.77±3.57 95.82±0.81 96.14±0.88 97.14±0.49 98.09±0.42
Figure 8: Visualization of the classification maps on the Houston dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-D Ablation Study
III-D1 Performance Contribution of Each Module
We explore the impact of each module within SFormer on classification performance through an ablation study, as shown in Table I. The baseline architecture utilizes a single $3\times 3$ convolutional layer, as well as a variant that employs nine consecutively stacked $3\times 3$ convolutional layers to align with the depth of the SFormer network. The results presented indicate that both the KSTB and TSTB modules significantly enhance the network’s classification performance. For example, on the Pavia University dataset, the baseline with single convolutional layer achieves an OA of only 87.67%. However, increasing the number of network layers did not lead to improved accuracy. In fact, employing nine convolutional layers yielded an OA of only 86.33%. The introduction of the KSTB module improved the OA to 93.44%, while the application of the TSTB module increased the OA to 95.81%. Each of these modules contributes to improved classification performance through their respective mechanisms. Specifically, the KSTB module improves recognition accuracy for objects of varying scales by enabling the network to adaptively select appropriate receptive field ranges from spatial and spectral selection. While the TSTB module enhances classification by selectively computing self-attention on the most relevant tokens with group strategy, effectively filtering and utilizing contextual information. This process reduces redundancy and interference from irrelevant information, thus improving the integration and utilization of useful information. When both KSTB and TSTB modules are applied together, the classification performance is further improved, with the OA yielding 96. 59%. In summary, the results of the ablation study demonstrate that KSTB and TSTB modules improve the network’s classification performance to varying degrees across all four HSI datasets, while the combination of two modules exhibits a synergistic effect, achieving the best OAs and confirming the effectiveness of these modules in enhancing network classification performance.
TABLE V: Quantitative Classification Results of the Indian Pines Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
1 93.87±2.26 93.23 ± 2.26 99.68 ± 0.97 90.32 ± 6.12 90.00±7.83 96.77±4.33 100 97.10±3.04 99.03±1.48 99.35±1.29
2 60.45±9.78 75.64 ± 1.76 89.33 ± 1.66 80.86 ± 6.15 68.72±3.71 84.45±7.24 88.08±3.69 88.15±2.65 89.47±2.85 93.22±2.14
3 70.64±4.44 83.36 ± 2.00 93.59 ± 1.95 94.21 ± 2.93 82.14±6.69 79.05±26.67 93.76±3.85 94.05±5.19 96.53±2.67 96.08±2.40
4 95.72±3.80 92.51 ± 3.08 97.86 ± 1.29 95.29 ± 1.24 94.60±2.40 97.65±2.74 99.63±0.79 99.89±0.32 99.89±0.21 99.79±0.43
5 89.49±2.49 90.30 ± 3.39 97.76 ± 0.75 94.83 ± 3.19 90.81±2.73 95.38±2.33 96.56±2.07 96.95±1.59 97.23±1.69 98.48±1.45
6 96.85±0.64 89.75 ± 1.65 98.65 ± 0.79 96.74 ± 2.12 95.32±1.80 96.82±5.89 99.34±0.42 98.63±1.28 99.43±0.61 99.68±0.22
7 96.92±3.77 98.46 ± 3.08 100 100 98.46±4.62 100 100 100 100 100
8 98.06±1.90 94.35 ± 1.40 98.62 ± 0.82 99.77 ± 0.26 97.27±2.36 99.58±0.55 99.86±0.28 99.74±0.29 99.93±0.21 100
9 100 100 100 100 98.00±6.00 100 100 100 100 100
10 76.80±6.72 75.42 ± 2.65 89.34 ± 2.62 93.67 ± 2.42 81.40±3.86 88.37±5.57 91.91±2.68 91.34±3.36 94.03±2.51 94.03±2.55
11 66.88±7.56 64.51 ± 3.09 91.06 ± 1.24 89.84 ± 5.82 73.53±4.42 60.80±25.84 82.13±4.48 87.42±4.85 89.50±3.21 90.25±1.53
12 82.71±7.42 80.41 ± 5.46 86.10 ± 1.31 91.20 ± 2.78 74.84±6.65 91.14±3.34 94.97±2.23 95.49±1.51 95.82±1.88 96.24±1.73
13 99.87±0.26 98.39 ± 1.09 99.61 ± 0.43 100 99.42±0.54 99.94±0.19 100 99.94±0.19 99.94±0.19 100
14 92.24±2.85 89.34 ± 1.18 96.02 ± 1.04 95.74 ± 1.77 92.10±2.82 96.29±2.00 95.62±1.83 97.42±1.54 97.77±0.88 98.52±0.66
15 94.85±4.19 90.65 ± 2.18 99.97 ± 0.09 99.52 ± 0.40 93.45±4.24 97.11±1.79 98.72±1.20 98.72±1.40 99.38±0.80 99.88±0.15
16 100 99.30 ± 1.07 100 99.07 ± 1.14 99.53±0.93 99.77±0.70 100 99.77±0.70 100 100
OA (%) 78.28±2.01 79.40±1.30 93.05±0.60 91.93±3.22 81.88±1.36 83.47±9.07 91.39±1.32 92.91±1.52 94.26±0.89 95.17±0.93
$\kappa$ (%) 75.30±2.26 76.68±1.45 92.04±0.68 90.78±3.66 79.43±1.51 81.43±9.94 90.17±1.48 91.89±1.73 93.43±1.01 94.47±1.07
AA (%) 88.46±1.28 88.48±1.09 96.10±0.35 95.06±1.91 89.35±1.10 92.70±3.71 96.29±0.45 96.54±0.82 97.37±0.39 97.85±0.44
Figure 9: Visualization of the classification maps on the Indian Pines dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-D2 Performance Contribution of Spatial-Spectral Mechanism
We further explored the impact of the spatial-spectral selection mechanism within KSTB on classification performance. As shown in Table II, classification performance decreases in the four datasets when using the single spatial or spectral selection branch individually. This highlights the importance of considering both spatial and spectral information for the accurate identification of ground objects in HSIs. It is essential to not only adaptively select the spatial receptive field but also to carefully choose the appropriate spectral information, as the required spectral contextual information varies significantly across different land cover types. For example, in the Pavia University dataset, the OA reaches 96.09% and 96. 01% respectively when using the single spatial or spectral selection mechanisms. In contrast, when both spatial and spectral selection mechanisms are applied simultaneously, the OA reaches 96.59%. In summary, this ablation study demonstrates that optimal classification performance is achieved when both spatial and spectral selection mechanisms are employed, thus validating the effectiveness of the proposed spatial-spectral selection mechanism.
TABLE VI: Quantitative Classification Results of the Whu-HongHu Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
1 90.78±2.65 92.34 ± 1.76 88.22 ± 1.86 89.94 ± 2.38 95.28±1.04 82.37±15.01 96.13±0.81 94.97±4.13 93.46±5.86 96.35±0.77
2 92.10±1.50 88.75 ± 1.34 89.69 ± 1.26 88.54 ± 5.40 90.01±2.12 69.44±31.34 96.23±1.47 95.61±2.81 95.07±4.31 96.84±1.23
3 80.27±7.44 81.16 ± 0.94 85.64 ± 1.10 85.93 ± 0.73 78.32±1.68 80.07±11.96 84.00±8.94 87.78±4.15 89.40±2.69 92.66±1.20
4 90.10±2.58 87.93 ± 1.28 91.50 ± 1.90 89.94 ± 5.18 93.20±2.40 84.41±11.17 95.46±0.70 93.31±2.80 95.56±5.62 97.09±1.01
5 85.49±3.52 84.99 ± 1.04 90.95 ± 0.75 93.46 ± 5.21 85.74±5.17 72.78±19.67 93.45±2.38 94.44±3.14 90.68±9.77 96.76±1.79
6 90.33±4.36 85.74 ± 1.27 86.64 ± 1.04 91.05 ± 2.85 90.41±2.92 66.43±29.85 94.12±2.52 94.68±2.32 93.32±2.90 94.95±1.93
7 66.40±3.44 61.18 ± 2.27 57.62 ± 2.61 76.95 ± 3.22 70.49±3.32 69.35±14.28 78.60±3.92 78.57±3.69 80.01±5.05 84.44±1.99
8 59.23±6.12 51.81 ± 2.87 81.22 ± 3.34 97.54 ± 1.79 63.43±3.11 64.45±12.07 80.02±4.75 84.93±7.33 83.12±3.24 92.69±2.15
9 95.45±1.16 95.04 ± 0.74 87.93 ± 1.76 92.61 ± 1.73 96.94±0.86 95.58±3.40 97.27±1.74 96.12±2.53 97.55±1.37 97.21±1.80
10 82.54±4.28 66.45 ± 2.86 59.82 ± 2.38 86.82 ± 2.46 61.54±5.74 81.11±8.85 89.57±2.24 88.08±5.46 82.34±11.69 93.54±1.42
11 68.36±5.18 56.68 ± 2.51 57.55 ± 3.51 80.69 ± 4.14 71.51±3.36 55.64±19.38 85.85±3.57 87.15±4.04 87.65±3.80 92.58±1.61
12 70.88±5.60 58.97 ± 2.42 78.58 ± 3.28 89.90 ± 3.60 70.96±4.17 56.33±20.32 79.35±5.21 82.53±5.54 81.50±9.10 88.52±2.96
13 66.08±7.33 54.73 ± 1.88 67.93 ± 2.28 86.36 ± 4.36 67.83±5.20 67.83±5.20 78.49±4.74 80.09±4.50 78.77±4.33 84.46±2.80
14 86.10±5.35 76.66 ± 2.60 78.39 ± 1.78 95.91 ± 2.47 85.54±3.17 85.43±8.74 94.55±2.26 95.28±3.66 92.92±5.11 97.54±0.91
15 94.50±1.95 93.32 ± 1.00 98.58 ± 0.83 99.35 ± 0.47 94.02±2.90 97.13±2.83 98.26±1.01 98.94±1.06 98.89±0.66 99.26±0.76
16 91.36±3.75 89.41 ± 1.53 90.74 ± 2.92 94.29 ± 2.02 89.36±3.51 91.33±3.20 94.82±1.69 94.05±4.07 98.31±1.48 97.26±1.37
17 91.41±3.83 89.81 ± 2.16 95.75 ± 1.78 95.14 ± 2.61 89.21±2.37 92.08±3.87 96.65±1.96 97.54±1.77 95.30±5.49 97.66±1.49
18 91.13±2.78 90.33 ± 2.19 94.03 ± 0.81 97.23 ± 1.67 90.66±3.82 64.21±25.62 95.68±2.52 98.21±1.22 96.14±2.73 98.80±0.90
19 89.90±2.60 85.11 ± 1.81 81.46 ± 1.40 86.18 ± 2.97 90.03±1.94 67.38±24.16 94.89±0.94 94.36±1.48 94.21±2.12 95.80±0.74
20 92.40±2.76 96.38 ± 0.78 91.49 ± 2.31 99.56 ± 0.16 94.45±2.00 75.31±22.05 95.69±2.37 96.27±2.90 97.42±1.69 98.90±0.52
21 83.04±9.61 87.71 ± 1.64 95.23 ± 2.22 99.69 ± 0.20 87.86±5.67 55.38±36.65 93.64±12.16 92.61±15.09 89.48±11.80 99.01±0.66
22 90.53±4.40 87.92 ± 1.82 94.77 ± 3.57 99.13 ± 0.66 91.54±3.50 88.84±9.76 95.71±2.00 97.01±2.54 97.14±1.80 97.95±1.12
OA (%) 85.16±1.18 81.26±0.70 84.12±1.13 89.10±2.83 86.36±1.05 77.17±9.41 91.65±0.85 91.19±1.52 91.79±2.32 94.59±0.50
$\kappa$ (%) 81.67±1.39 76.98±0.81 80.33±1.31 86.55±3.32 83.03±1.24 72.46±10.69 89.55±1.05 89.04±1.82 89.76±2.72 93.20±0.62
AA (%) 84.02±1.09 80.11±0.72 83.81±0.68 91.65±1.31 84.01±0.83 75.03±9.24 91.29±0.90 91.93±1.60 91.28±1.23 95.01±0.21
Figure 10: Visualization of the classification maps on the Whu-HongHu dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
Figure 11: Comparison of classification performance with different training sample per class: (a) Pavia University dataset. (b) Houston dataset. (c) Indian Pines dataset. (d) Whu-HongHu dataset.
TABLE VII: Comparison of Model Parameters (K), Inference Time (s) and OA (%) Across Four Datasets
Dataset Method SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
Pavia University Parameters (K) 174.19 484.42 927.11 149.75 984.33 571.88
Inference time (s) 44.03 8.52 15.86 26.17 14.06 39.40
OA (%) 76.46 75.79 90.69 90.62 93.40 96.59
Houston Parameters (K) 235.95 673.74 973.58 155.19 995.22 577.90
Inference time (s) 226.92 37.94 60.49 99.22 123.07 136.28
OA (%) 88.10 86.72 94.99 95.36 96.52 97.66
Indian Pines Parameters (K) 352.44 931.85 830.80 199.99 1009.62 585.20
Inference time (s) 14.05 1.44 2.46 3.55 1.96 4.66
OA (%) 81.88 83.47 91.39 92.91 94.26 95.17
Whu-HongHu Parameters (K) 551.28 1254.80 1514.97 255.99 1027.93 594.93
Inference time (s) 274.36 45.98 63.54 83.65 49.25 108.51
OA (%) 86.36 77.17 91.65 91.19 91.79 94.59
III-E Performance Comparison and Analysis
In this subsection, we compare the proposed method with other classical and state-of-the-art approaches. A brief introduction to each method included in the experiments is provided below.
  • 1)
    3D-CNN[29]: Spatial-spectral classification using 3D-CNN. We follow the implementation with the original article.
  • 2)
    SSFCN[26]: Spatial-spectral classification with an end-to-end fully convolutional network. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a momentum of 0.9.
  • 3)
    SACNet[38]: Spatial-spectral classification method that integrates self-attention learning with context encoding. The network is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a weight decay of $55e-$ 5.
  • 4)
    FcontNet[36]: Multiscale classification based on non-local contextual attention across space, bands, and scales. The base learning rate is 0.01, adjusted using the poly scheduling policy. The model is optimized using SGD with a momentum of 0.9 and a weight decay of $41e-$ 4.
  • 5)
    SpectralFormer[22]: A Transformer-based architecture that employs sequential learning, focusing on adjacent bands and patches. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4, which decays by a factor of 0.9 after every tenth of the total epochs.
  • 6)
    SSFTT[43]: Spatial-spectral classification method integrates a CNN with a single-layer transformer and a Gaussian distribution-weighted tokenization module. The model is trained using the Adam optimizer with a learning rate of $31e-$ 3.
  • 7)
    GAHT[48]: Spatial-spectral classification through a hierarchical transformer with grouped pixel embedding. The model is trained using the SGD optimizer with a momentum of 0.9, a weight decay of $41e-$ 4, and a learning rate of $31e-$ 3.
  • 8)
    MorphFormer[49]: A ViT-based spatial-spectral classification method that combines attention mechanisms with morphological operations. The model is trained using the Adam optimizer with a weight decay of $35e-$ 3 and a learning rate of $45e-$ 4.
  • 9)
    GSCViT[45]: A lightweight network utilizing groupwise separable convolution and ViT in each feature extraction block. The model is trained with the AdamW optimizer, featuring a learning rate of $31e-$ 3 and a weight decay coefficient of 0.05.
Figure 12: Visualization of the objects with different receptive fields selection weights from Pavia University dataset. (a) Position of three objects P1: Asphalt, P2: Meadows, and P3:Bare soil. (b) Weights selection values of different receptive fields.
III-E1 Quantitative and Qualitative Results Analysis
The quantitative results include the accuracy per class, OA, $\kappa$, and AA as shown in Table III to Table VI, with the best results highlighted in bold. Taking the Pavia University dataset as an example, CNN-based HSI classification methods such as 3D-CNN at the patch level, the classification performance is limited by the size of the input patch, resulting in an OA of only 77.01%. For image-based methods such as SSFCN, although the entire image is used as input, the limited receptive field of the convolutional layer hampers the effective extraction of contextual information, leading to a lower OA of 81.48%. The SACNet method introduces a self-attention learning mechanism and context encoding to establish long-range dependencies, but its performance improvement is restricted by spatial resolution during the self-attention mechanism calculation, yielding an OA of 83.27%. In contrast, FcontNet achieves a relatively higher accuracy with an OA of 90.54% by combining scale, spatial, and channel attention mechanisms for multiscale classification. However, the extensive self-attention computations and multiscale feature design of this method result in high computational resource demands and low operational efficiency.
For transformer-based methods, both SpectralFormer and SSFTT are built on the ViT structure and learn by constructing sequence data from adjacent bands. However, due to redundancy between bands and the lack of spatial context modeling, these methods perform slightly worse than the CNN-based HSI classification methods aforementioned, with an OA of 76. 46% and 75.79%, respectively. GAHT builds a hierarchical transformer on the ViT foundation by integrating grouped pixel embeddings, MorphFormer incorporates morphological operations into the ViT attention mechanism, and GSCViT uses groupwise separable convolution with ViT to extract local and global features. All of these methods achieve relatively high accuracy, with OAs of 90.69%, 90.6%, and 93.40%, respectively. However, these approaches overlook the varying range of background information required to identify different types of land cover in the HSI, and the traditional self-attention mechanism includes redundant and irrelevant information during computation. Furthermore, insufficient refinement in capturing the interspectral context limits the accuracy of existing transformer-based HSI classification methods.
Figure 13: Visualization of (a) the dense attention matrix and (b) selective attention mask in Pavia University dataset.
To address the aforementioned issues, we employ KSTB for adaptive receptive field selection, followed by TSTB to intelligently choose the most relevant tokens for self-attention computation. Our method, SFormer, achieved an optimal OA of 96.59%, surpassing the second-best method by 3.2 percentage points. Notably, SFormer obtained the best results in nearly all per-class accuracy metrics, except for the ninth class, shadows, where it was slightly outperformed by MorphFormer. This may be attributed to the integration of morphological operations within the attention mechanism, which likely provides more effective semantic features for shadow recognition. Similarly, our proposed method demonstrated superior accuracy in three additional datasets, highlighting its advantages and applicability.
We also present classification maps in Fig. 7-10 to evaluate various methods. Using the WHU-HongHu dataset as a case study, we observed that methods such as 3D-CNN and SSFCN exhibit significant salt-and-pepper noise. In contrast, FcontNet and the latest Transformer-based approaches yield smoother and better results. A particularly noteworthy observation is on the classification of the cotton class, where most existing methods tend to misclassify it as cotton firewood. However, our SFormer model significantly reduces these errors, producing outputs that are much closer to the ground truth. This advantage has been consistently validated across multiple datasets, highlighting the superior performance of the SFormer model in handling complex classification tasks.
III-E2 Robustness with different training samples
To assess the stability of the proposed SFormer method, we conducted extensive experiments on four diverse datasets, each with varying numbers of training samples. Specifically, we randomly selected 25, 50, 75, and 100 samples per class to construct the training sets. Fig. 11 compares the classification performance in these different sample sizes. The results indicate that the classification accuracy of SFormer is steadily improving as the number of training samples increases, demonstrating its robustness. Furthermore, SFormer consistently outperforms existing state-of-the-art methods, achieving the highest accuracy across all training sample configurations.
III-E3 Computation Cost Analysis
Given that previous comparative experiments indicate that transformer-based models typically achieve higher accuracy, we performed a comprehensive comparison of all transformer-based methods regarding parameters, inference time, and OA results, as presented in Table VII. In our performance evaluation, the proposed SFormer exhibited strong competitiveness and achieved a good balance between precision and efficiency in four benchmark datasets. Regarding model parameters, our SFormer has fewer parameters than the transformer-based state-of-the-art methods GAHT and GSCViT. Although the inference time of SFormer is slightly higher than that of the comparison methods, it outperforms SpectralFormer. These results suggest that SFormer successfully achieves an optimal trade-off between computational resource usage and the ability to capture complex features.
III-F Visualization
III-F1 Different Objects of Selection Weights
To validate differences in receptive field requirements for various types of land cover in the HSI classification, we present the visualization of objects with different selection weights for receptive fields. Fig. 12 (a) presents the locations of three land covers: Asphalt, Meadows and Bare Soil within the Pavia University dataset and their corresponding positions. And Fig. 12 (b) shows the weight values $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ for different receptive fields used in the final fusion within the KSTB. The weight results show that the weight corresponding to the larger receptive field in P1 is 54. 07%, which confirms our previous analysis that asphalt requires more contextual information to perform an accurate classification. In contrast, at locations P2 and P3, both of which exhibit spatial continuity, smaller receptive field characteristics are sufficient for classification needs.
III-F2 Attention Matrix Change of Token Selection
To further understand the effect of the token selection operation, we present the changes in the attention matrix before and after the operation. As shown in Fig. 13, we used the Pavia University dataset with the number of groups set to 2 and a selection rate of 40%. Specifically, we selected the dense attention matrix from the first head of the multi-head self-attention mechanism. From the figure, it is evident that after the token selection operation, only 40% of the elements in the original dense attention matrix retain their values, while the remaining 60% are set to 0, resulting in a sparse selective attention mask. After deep feature extraction by the TSTB, the relevant information for classification is retained, whereas irrelevant and redundant information is discarded.
IV Conclusions
In this article, we introduce a novel SFormer for the HSI classification. The SFormer is designed to dynamically select receptive fields to capture both spatial and spectral contextual information, while reducing the influence of redundant data by emphasizing the most relevant features. Recognizing that objects in HSI of varying scales require distinct spatial-spectral contextual information, the SFormer incorporates a KSTB, which dynamically selects appropriate receptive fields for effectively spatial-spectral feature extraction. Furthermore, considering vanilla self-attention mechanisms may be susceptible to redundant and irrelevant information, which can negatively affect classification performance. We introduce a TSTB that selects the most valuable tokens with the highest attention values, rather than applying dense attention across all query-key pairs. These selective mechanisms allow SFormer to capture relevant information within the optimal receptive field, enhancing its ability to effectively extract valuable spatial-spectral features. Extensive experiments conducted on four benchmark HSI datasets demonstrate that the proposed SFormer achieves competitive performance compared to existing state-of-the-art methods, thereby validating the effectiveness of the proposed SFormer.